### v2

# Storypark: Leveraging Large Language Models to Enhance Children Story Learning Through Child-AI collaboration Storytelling

Abstract.
Interactive storytelling has been widely adopted by educators in teaching activities of young children. Such a teaching method combines storytelling with active child participation, benefiting their expressive abilities, creative thinking, and understanding of stories. Interactive storytelling requires facilitators to unidirectionally narrate the story content and encourage children’s participation in story plot creation and interpretation of central themes through multi-sensory interactive methods such as questioning and drawing. However, providing tailored guidance based on diverse feedback from children during interactive storytelling poses challenges for most facilitators. These challenges include expanding story plot development based on children’s ideas, using drawings to visualize children’s thoughts, and interpreting the story’s central themes based on children’s thinking. This necessitates facilitators to possess strong imaginative, associative, domain knowledge, and drawing skills. Large language models have demonstrated their potential in facilitating responsive and participatory dialogues, offering new design possibilities to address the challenges faced by facilitators in interactive storytelling.
In this study, our goal is to leverage large language models to design an interactive storytelling system that provides children with plot frameworks and interpretations of central themes during the interactive storytelling process. Through user experiments involving 20 child participants, we evaluate this interactive system’s usability, learning effectiveness, and user experience. The user study shows that Storypark improves learning outcomes in understanding story key ideas, generalization, and transfer. And high engagement and willingness to use of participants demonstrate that StoryPark provides children with a positive learning experience.
Human-centered computing Interactive systems and tools; Storytelling, Children, Multimodal interfaces; Large language models; Child–AI collaboration, Co-creativity
$c$s: Human-centered computingccs: Human computer interaction (HCI)$c$s: Interactive systems and toolsccs: User interface programming
1. Introduction
To enrich children’s learning experiences before daily conversations and scenarios, researchers emphasize the importance of incorporating storytelling materials into teaching activities. Story materials can help children to have early exposure to domain knowledge and some universal principles of human society in diverse and dramatic scenarios. Interactive storytelling is a flexible and effective educational method that integrates storytelling with children’s participation, and it has been widely applied in teaching activities for preschool and elementary school children. This educational approach requires facilitators to center their interaction around story materials, engaging with children through various forms. Among these, the most common form of interaction is interactive dialogue. In interactive dialogue, facilitators pose questions to children about the story or pictures in the book and provide scaffolded feedback based on children’s responses. Feedback strategies include implicit correction, repetition, expansion, and modeling of answers in various forms. Within dialogic reading, facilitators encourage children to participate in the creation of story plots by designing questioning methods and articulating their understanding of the core ideas contained within these storylines. Compared to one-way storytelling, interactive storytelling does not cultivate passive listening; instead, this process enhances children’s understanding of story meaning, language expression abilities, and creative thinking. Previous research has found that increasing interactivity and semantic guidance during storytelling can effectively enhance the positive impact of storytelling on children’s language development. Beyond language, facilitators also encourage children to visualize their ideas through drawing.
Therefore, this story-based learning method requires children to comprehend and practice the educational content of stories under the guidance of more knowledgeable individuals such as parents, teachers, or older peers. However, providing real-time guidance to facilitators based on children’s feedback during interactive storytelling poses numerous challenges. These challenges include expanding new storylines based on children’s ideas and integrating them with the story background, visualizing children’s ideas through drawing, and further explaining the core ideas of the story based on children’s understanding. This necessitates facilitators to possess strong language expression abilities, imagination, associative thinking, broad domain knowledge, and preferably, some level of drawing skills. Additionally, interactive storytelling necessitates a one-on-one environment to ensure children’s active participation. However, given the limited resources faced by most educational institutions, one-on-one storytelling educational resources are not feasible in reality.
To address this challenge, researchers have long been exploring the potential of artificial intelligence (AI) to simulate human storytelling interaction. Prior to the emergence of large language models, AI primarily focused on optimizing structured interactions with children, such as questioning and providing evaluative feedback. With advancements in AI technology, large language models have demonstrated their potential in facilitating responsive and participatory dialogue, generating contextually relevant and coherent text based on the user input received so far. This capability can simulate ”real dialogue,” offering new design possibilities for addressing the challenges faced by facilitators in interactive storytelling. However, the feasibility and effectiveness of involving children in storytelling through large language models as partners remain unclear, especially in the context of supporting contextual understanding learning. Furthermore, it is not yet clear what specific workflows are needed to customize the capabilities of large language models and design interactive systems for storytelling collaboration between children and AI, tailored to the needs of child users in interactive storytelling scenarios.
In summary, this paper makes the following contributions:
  • Design and develop Storypark, a multi-modal child-story creation system that enhance story learning including core value and domain knowledge through the use of a question-scaffolding feedback framework. Provide multi-modal creativity support for children by conversational generation assistant and sketch generation assistant.
  • Formulat the prompt engineering of children-story cocreation for educational purposes base on large language model. We evaluate the performance of artificial intelligence compared to human users in generating expanded story content and questions for the task of storytelling to children.
  • Results from a user study, summarized the advantages and limitation of Storypark from the aspects of learning effectiveness and learning experience, as well as the qualitative views of children, parents and education experts on Storypark.
2. Related works
In this section, we review research on: (1) Teaching values and knowledge through interactive storytelling, (2) Multimodal collaboration system, and (3) Using LLMs for child-facing storytelling system.
2.1. Understanding values and knowledge through interactive storytelling
Interactive storytelling has been demonstrate as an effective educational method(Gallagher, 2011; Collins, 1999; Bonds, 2016; Leonard, 1990). And it has been proven that it can improve children’s language and narrative skills(Anggryadi, 2014; Zuhriyah, 2017). In the human-computer interaction community, there have also been some works in natural language processing focusing on issues related to interactive storytelling(Zhang et al., 2022; Yao et al., 2022; Marzuki et al., 2016; Xu et al., 2022). However, these studies focus on how to improve the performance of interactive storytelling system, but do not pay attention to the learning process of children’s interaction with the story system. With the emergence of large language models, some studies have incorporated large language models into children’s interactive storytelling activities. Nisha use automated planning in Natural Language text generation to create stories that are coherent and believable(Simon and Muise, 2022; Simon, 2024), jiajun introduce the FairytaleCQA dataset for children’s story narritive. but these study have not take into account how the system interacts with children. yuling sun’s qualitative research proves that parents have a need for interactive story systems, but her investigation focuses more on tangible interfaces and has yet to come up with a solution(Sun et al., 2024). In addition, studies have explored the use of llm to promote children’s cooperative activities (jiawen liu, 2023a; Zha et al., 2024), as well as parent-child interaction (Lee et al., 2022), child socializing (Seo et al., 2023). However, these studies focus on small areas of application, and interactive storytelling is a basic and general educational tool that can have a wider range of applications.
2.2. Multimodal collaboration system
Previous studies have designed story idea generation tools for children. Generation methods include vr avatars, voice interface creation stories(Barber and Kudenko, 2010), through drawing creation, tangible interface(Liang et al., 2021) and graphic interface(Shao et al., 2021; Murray et al., 2017). Such tools underscore the significance of multi-modal creative support, enabling children to express themselves through various mediums. However, few studies have focused on addressing the core truth of children’s understanding of stories through participatory creation. Benton present a storytelling system for narrative co-creation through a game(Benton et al., 2014), Zarei came up a enactment-scaffolded narrative authoring tools(et al., 2021). But these studies do not integrate the educational goals and understanding processes of stories. Some studies have noted that interactive systems improve children’s math(Zhang et al., 2024; Ruan et al., 2020) , coding skill(Rocha et al., 2023; jiawen liu, 2023b) and narrative skills(Liang et al., 2015). Some studies have noted that interactive systems improve children’s math and narrative skills. But most of the research has tried to improve a single skill. Few studies have focused on designing an interactive system that helps children improve the more general skills of understanding and transferring the core values of stories through interactive storytelling.
2.3. Using LLMs for child-facing storytelling system
Pretraining large language models, such as GPT-3(Brown et al., 2020), GPT-4(OpenAI, 2023), and LLAMA-2(Meta, 2023), have accelerated the field of NLP in recent years. These models are trained on large amounts of text data and can perform downstream tasks such as text generation(Wei et al., 2022) etc. Since the large model output depends on the input, the prompt of the large model greatly affects the output performance of the large model, and the prompt allows individuals to provide specific text instructions and examples to guide the model to perform the task(Zamfirescu-Pereira et al., 2023). Although the nudged llm may not consistently outperform baseline models, it provides a lightweight approach that achieves competitive performance on a variety of tasks(Brown et al., 2020; Chowdhery et al., 2023).
LLMS have the potential to facilitate responsive and participatory conversations. They are not limited by a predetermined dialogue paradigm, nor do they rely on pre-written replies. LLMS are able to generate contextually relevant and coherent responses based on the input they receive(Wei et al., 2022), which allows them to achieve near-real conversations, collaborate effectively with children, and create a cohesive response. Generate naturally rich dialogue(Lee et al., 2023a). Therefore, llm is very suitable for the field of education and interaction with children. Many researchers have made efforts to incorporate llm into educational technology(Kasneci et al., 2023). These adaptations witness the application of llm in creating educational content(Dijkstra et al., 2022; Gabajiwala et al., 2022), enhancing student engagement and interaction(Abdelghani et al., 2023; Bao, 2019; Tai and Chen, 2023) and personalized learning experiences(Sailer et al., 2023). For example, LLMS have been used to generate children’s narratives(Alabdulkarim et al., 2021; Guan et al., 2020). In different research trajectories, some scholars have used LLMS to create intelligent learning partners that are able to collaborate with humans(Ji et al., 2023a), provide feedback(Jia et al., 2021) and encourage students(El Shazly, 2021; Tai and Chen, 2023). A common application involves employing LLMS as dialogue partners in written or oral form, such as in the context of task-oriented dialogues that provide opportunities for language practice(El Shazly, 2021). Or use llm as a story continuation partner to collaborate with children to create stories(Peng et al., 2023). Overall, we believe that large language model applications, can provide children with unique conversational experiences that complement their interactions with others. Our findings have implications for designers and developers of applications based on large language models designed to support children’s learning. In addition, they highlight areas that require further consideration to ensure that the technology is both supportive and effective for different learning experiences.
3. Storypark: System design
To generate the design principles of the participatory story interactive system for children, we recruited five education experts and three parents to conduct semi-structured interviews. Each participant had more than one year of experience in educating or raising children. The interview focused on three points: the educational purpose, the benefits of interactive storytelling for children, and the problems that educators currently encounter in interactive storytelling.
Participants indicate that telling stories to children generally has teaching purposes. Teaching purposes are usually divided into two types, daily domain knowledge and core ideas. Interactive storytelling aimed at children places high demands on the guide’s own knowledge, imagination, and narrative skills. Children, who have not yet undergone extensive learning processes, often provide feedback that can be unexpectedly surprising to adults. The guide needs to assist the children in weaving their imaginative responses into the story framework, elaborating new plot lines based on the children’s feedback. In the model evaluation mentioned below, we designed a questionnaire with a given story outline and educational objectives. Participants were required to devise interactive questions and write plot expansions and moral interpretations based on presumed children’s responses. Of the 20 participants, 15 had professional experience in early childhood education, while 5 had experience raising children aged 3-8. The participants commonly expressed that ”the questionnaire was challenging to answer” and ”without having been a kindergarten teacher, one would probably not manage to respond.” The questionnaire consisted of 9 questions, with the median time spent filling it out being 1638 seconds. This process also demonstrates the challenges that guides face in co-creating stories with children.
We followed these 3 design guidelines:
  • DG1: The storytelling system should adopt an interactive model of ”Question-Feedback-Story Generation”, presenting open-ended questions to elicit narrative contributions from children. Based on children’s contributions and the story framework, the system should expand details to support further story development.
  • DG2: The storytelling system should help children understand the core morals of the story and provide detailed explanations based on children’s responses and story samples.
  • DG3: The storytelling system should augment the ability for children to create stories through voice by incorporating drawing features for visual completion, providing additional visual cues for children.
3.1. Feature design
Figure 1. Child interaction process
Story can be divided into four parts includes domain knowledge, plot development, main idea, and story elements. Among them, knowledge and key values are the most important assessment points for learning stories, and they are also the teaching purposes that parents and teachers want to achieve through stories. The co-creation process between AI and children focuses on story development. We will describe the specific design details of the company park below. StoryPark uses AI capabilities to help children expand the contributions they want to add to the story, and effectively weaves the key values and knowledge points of the story into the development of the storyline. StoryPark has five interactive steps in this interactive storytelling process: 1) generate question leads to elicit children’s narrative contributions, and 2) expand storyline development based on children’s feedback input. 3) Help children complete story elements after they draw them. 4) Based on the children’s participation in the creation of the story, generate the truth questions to evaluate the central idea of the story. 5) Agree with the child’s answer and further explain the core truth expressed in the story. Figure1 shows the overview of the StoryPark process.
3.1.1. Question-Feedback-Story Generation model
A round of dialogue sets up the story and listens to the background of the story. The AI asks the child questions and lets the child decide what happens at this point in the story. Two questions are designed, and the number is determined according to the suggestions of parents and preschool teachers. Because children’s attention is not able to focus for long periods of time, two questions can fit into the attention span of most children. At the end of the story, ask the child questions and explain the reason again. Multiple rounds of dialogue can take place here, and if the child gives an answer contrary to the key values of the story, the AI will give advice and give the right guidance. Two questions are designed to address the different core truths a story can convey.
3.1.2. Visual-voice interface: Multi-modal assistant
In the process of education, it is important to give children specific visual descriptions in time to attract children’s attention. It can help children visualize their imagination. Previously, educators had also used the method of drawing sticks on the blackboard. Therefore, we designed the function of AI to complete children’s drawing. The child first uses the button to record voice, and then uses the sketching board to directly draw the story they want to create. In the migration scenario, StoryPark randomly provides story elements created in previous stories, as well as elements from other stories. The voice tells the child to use different story elements to convey the same core idea. Test whether children can combine the truth they have just understood with the domain knowledge they have previously understood to build new stories. The child needs to transform different scenes and story elements and be able to describe the development of the story using sounds and images respectively.
3.2. How to use this system
As shown in the figure2, Storypark is used on iPad. Users press the voice button to communicate with Storypark. When the voice button lights up, it means recording is in progress, and when the child presses the speak button again, the recording ends. Storypark will continue the story development based on what they have said, while still conveying the original core key value of the story. After the child tells Storypark about their creation, they can press the drawing button to bring up the drawing board, where children can freely draw elements of their creation directly in the story scene. Storypark has designed a feature to help children complete drawing elements because concrete drawing operations and visual feedback are helpful for children to increase their engagement with the story, deepen their understanding, and aid in memory retention. After completing the entire story, Storypark will ask the children two questions in succession to assess whether they understand the knowledge and core key value of the story. Children can engage in multiple rounds of voice conversations with Storypark. Even if children provide incorrect, negative (”I don’t like being polite”), irrelevant, or uninformative answers, Storypark will patiently explain the core knowledge of the story to them. In the final transfer phase of the assessment, children are asked to use Apple Pencil 3 to draw or erase on the canvas and switch colors in the drawing tool area. The system also allows children to move, rotate, or scale the story elements provided on the canvas through touch gestures such as dragging, rotating, and pinching. Based on these elements on the canvas and the key value of the story just co-created, they can create a new story that still conveys the same key value.
Figure 2. User Interface of StoryPark
Here is a complete interactive process example:
  1. (1)
    StoryPark[prologue]: Once upon a time, there was a little turtle…
  2. (2)
    StoryPark[first interactive question]: Children, what can bear do with its shell?
  3. (3)
    Children[pressing the voice button]: I know bears like honey and can make a bowl out of shells.
  4. (4)
    Children[Press the voice button again to indicate the end of this audio recording]
  5. (5)
    Children[press sketch button], StoryPark[evoke sketching board], Children[start sketching]
  6. (6)
    StoryPark[SketchRNN helps children complete drawing details based on classes choosen by children]
  7. (7)
    Children[Press the sketching button again to represent the end of sketching], StoryPark[Collect elements drawn in the sketching board]
  8. (8)
    StoryPark: Wow, kid, you’re so smart! Like you said, little Bear uses the turtle’s shell as a bowl…
  9. (9)
    StoryPark[second interactive question]:… Omit the second round of interactive questions and the second round of story creation
  10. (10)
    StoryPark[end of story]: You’re awesome, kid! The bird uses the turtle’s shell as… This process changes the way the turtle thinks about its shell…
  11. (11)
    StoryPark[first key value question]: Kids, why does the bird thank the turtle?
  12. (12)
    Children[presses voice button]:” Shell is good”
  13. (13)
    Children[Press the voice button again to indicate the end of this audio recording]
  14. (14)
    StoryPark[first key value guide]: The bird thanks the little turtle because… What this tells us is that sometimes even seemingly ordinary traits or things…
  15. (15)
    StoryPark[end of story, random display of the elements the child has just created, as well as other story elements]: Kids, you can use these elements to tell a new story, still express…
4. Implementation
In the concrete implementation, StoryPark consists of two parts, namely prompt engineering and system building. Since LLMs is used as the core component of the multimodal interaction process, the prompts used to utilize LLMs execution are analyzed and designed in detail to ensure stable LLMs performance in all aspects. For the part of system building, we first evaluated the current LLMs and human performance on continue writing and feedback, verified the feasibility of using LLMs. Secondly, we explained the technical implementation of the system in detail, and showed the technical implementation details of StoryPark.
4.1. Prompt Engineering
4.1.1. Question Generation
In order to achieve interactive storytelling and meet the requirement of asking open-ended questions in the design principles described, the design of questions in the story should avoid simple Y/N questions, allow children to make diverse answers, and the direction of the answer should not exceed the original frame of the story. Therefore, we will send a preset prompt to the model before question generation, such as ”Tell a story to a 3-7 year old child, and let them understand the positive value of the story, given a single open-ended question at the end of the unfinished story, which answer can expand story” In general, depending on the story frame, we directed the model questioning about item usage or language response. At the end of the story, StoryPark will raise questions about the key value contained in the story. We set prompt let model to generate questions with positive educational effect and in line with the story outline, and give positive feedback after children answer. For possible deviations from educational purpose or negative answers, the model will give guidance in the correct educational sense and give encouragement.
4.1.2. Story Continuation
Role setting for large language models: In order to better enable LLM’s to continue writing interactive stories, and ensure the establishment of the overall story interaction framework, we first prompting the model to assume the role of a ”interactive story telling robot”. At the beginning of each interaction process, we set the character prompt ”Your character is a children’s story interactive robot, and needs to interact and continue the story based on children’s feedback, and can convey positive values. Your audience is 5-12 years old.”
Story opening template Settings: The next part of the prompt focuses on the story context setting. We provide the model with a narrative scenario: ”The main outline of the story and the plot development is [story setting], and through this story teaching the child ’[key value]’. Please expand and continue this story, in the story, design interactive questions 1 and 2 ”. The square brackets [story setting] and [key value] represent the outline of the story development and the educational key value of the story, respectively, where the elements will be dynamically filled according to the story selection. We also set some high-level rules to ensure that the generated story content is ”simple and appropriate for young children” and does not contain anything unrelated to the story.
Story continuation template Settings: In order for the model to have coherence in story continuation before and after interactive questions, we need to indicate the model continuation requirements. In the continuation phase, the task of the model is to encourage children’s answers and give positive feedback, and then continue the content according to the story outline combined with children’s feedback. This is done with the following detailed prompts: ”Tell the beginning of the story, then continue the story to the position of question 1 and stop to generate interactive question 1, where you need to stop and wait for the child’s feedback”, ”The child’s feedback is [feedback], generate an expansion and continue the story until the child stops at question 2 and waits for the child’s feedback”, ”The child’s feedback is [feedback], generate the ending plot”, and the square brackets [feedback] represent the child’s feedback content.
Context maintenance in story continuations: We observed the LLM’s occasionally forgot the storyline which has already expanded as interactions continued and story outline of the setting. This results in LLM’s expands a completely unrelated new story based on the last answer. Since the OpenAI API interface is stateless, we maintain a dialogue context during the interaction, which saves the LLM’s dialogue with the child and the system prompt setting in the current round of interaction. This ensures the consistency of Storypark in story development.
4.1.3. Scaffolding
In the process of interactive story creation, children may give feedback that they cannot expand the story. We classified the situations where children require scaffolding into two main categories :(1) cases where they completely cannot continue the story, expressed as uncertainty (e.g., saying ”I don’t know”); (2) cases where they gave only short and vague answers, and the image completion assistant models did not include children’s answer object. To deal with this situation, we give different scaffolding, and to construct the first situation, our first step is using model to determine that the child is showing signs of uncertainty. After the uncertain answer is given, we utilize model to provide hints related to the question, and adhering ”paraphrasing” strategy. When solving the second case, GPT-4 model was used to comprehensively consider and feedback according to the children’s previous short answers and the list of image completion model types of image completion assistant. This method involves first acknowledging what the child said, continuing the story considering both related object in the image completion models and child’s feedback.
4.1.4. Ethical Considerations
The safe use of AI by children has been a topic of public discussion. As we transition to an era dominated by generative AI, the debate becomes more urgent. In this new context, children may directly encounter content produced by AI without the protective filters traditionally provided by human supervision.
On the privacy side, we utilize two cloud-based services to process children’s utterances: iFlytek’s speech-to-text for transcription, and OpenAI’s GPT-4 for story continuation and feedback generation. During the invocation, the child’s utterance is not stored. Additionally, OpenAI promises not to use data passed through the API to train their models. But when cloud-based AI products enter the market, the adequacy of children’s data privacy protection is still uncertain due to the lack of specific regulations for AI products. On the security side, to prevent inappropriate AI-generated content, we added a moderation layer that exploits OpenAI APIs to ensure the output is does not contain sexual content, hate speech, harassment, violence, or self-harm. We also use prompt Settings to limit the amount of content involved in story continuation and problem feedback.
4.2. Multi-modal Story Interaction System Build
StoryPark is implemented as a web-accessible software to run on an iPad with a iPencil, microphone, and speaker. The application is written in Python, JavaScript and Html to interact with stories through voice input, iPencil drawing, and tap operations.
The front end of StoryPark is built using the React framework, which is an open source JavaScript library maintained by FaceBook which uses components to quickly develop single-page applications. The sketch drawing components and image related operations the interface are implemented using the p5.js library. When sketching, the child can perform actions such as selecting brush colors, clearing the sketch board, completing the image, etc. The back-end server is developed using Flask framework. In the process of interaction, the data types of the front end and back end interaction include audio, text and image, so the StoryPark back end realizes the transceiver interface of multiple data types. We chose to use GPT-4 as our LLM model of choice, which is currently the most powerful LLM available from OpenAI, We use the prompt designed in 4.1 to send and receive the generated text content using OpenAI’s GPT-4 interface.
Figure 3. Multimodal interaction process technical details
Figure3 shows the multi-modal interaction process technical details. In multi-modal interaction, StoryPark implements the sketch module to provide the ability of sketch completion. Considering the differences in drawing ability of children of different ages, StoryPark will supplement the sketch after children have finished drawing. sketch-RNN is a recurrent neural network (RNN) capable of building common object graphs based on strokes using the dataset Quick Draw! trained on, the Quick Draw dataset is a collection of 50 million drawings across 345 categories. In the interaction process, firstly, the sketch module extracts the keywords in the child’s voice response and matching with 114 RNN model categories that realize the completion ability. The similarity matching is realized through LLM’s ability. After obtaining the keywords, the corresponding type of RNN model is loaded for sketch completion.
Through the voice module, StoryPark enables converse capabilities with children. In the voice narration process, after obtaining the generated text through prompt engineering, we use the text-to-speech interface provided by iFlytek to create corresponding voice audio, which is played to the children via the front end. In voice recognition, after the child presses the dialogue button, the iPad activates its built-in microphone to start recording. After the child clicks the dialogue button again, an audio file is created and sent to the backend voice module. We use iFlytek’s speech-to-text interface to recognize the content of the speech, and the recognized text content is inserted into the prompt to continue story generation.
4.3. Model Performance Evaluation
To assess the performance of GPT-4 in generating stories expansions and questions, we design a set of identical story outlines and educational points, which are distributed to both human participants and ChatGPT. By comparing the content generated by humans and ChatGPT, we evaluate the GPT-4’s performance in co-creating story plots, formulating open-ended core idea questions, expanding storylines, and providing detailed explanations of the story’s central concepts.
Figure 4. The Questionnaire of Human-generated story materials
We gathered 20 participants and generated 80 stories. Of the participants, 15 had professional learning experience in early childhood education and 5 had parenting experience with children aged 3-12. Participants were asked to write interactive questions. The author observed children’s reaction in these story nodes through a brief pre-experiment, and randomly selected some children’s feedback into the questionnaire. The questionnaire listed the designed questions and randomly selected three different children’s feedback, and participants need to expand the plot and explain the reason according to the questions and the three kinds of children’s feedback. As shown in Figure 4, for the four questionnaires issued, I added story pictures to facilitate participants’ understanding.
The questionnaire question design for human participants was the same as the prompt word task for ChatGPT4, and the same questions and guidance were used in storypark’s story core co-creation process. As Figure 4, the human participant and ChatGPT4 are required to complete the same task.After obtaining 80 stories from participants with early childhood education experience, we also generated 80 stories using GPT-4.
4.3.1. Evaluation index design
As shown in Figure 5, we designed a complete evaluation metric for the requirements of story co-creation. It is divided into two modules: text generation and question generation. The text generation module is divided into plot extension text and truth extension text, and the question generation module is divided into plot question and truth question. There are similar indexes and different indexes of the two groups of modules. The metrics are based on a widely accepted approach to evaluating AI-generated conversations presented in the work of Junyi Li and Ziwei Ji(Shi et al., 2023; Lovato and Piper, 2019). In this method, evaluators were shown pairs of stories, one generated by GPT-4 and one generated by a human. They were asked to choose from each pair based on the evaluation dimension. Therefore, we ask three evaluators to choose from each pair of generated text based on the different dimensions given in the index system. All assessments were conducted by three university student research assistants with more than one year of experience in the education of children.
Figure 5. Evaluation index design
4.3.2. Evaluation of Content generation
In the text generation evaluation module, the four evaluation dimensions of story generation are understandability, preference, logical rationality and attractiveness. For guidance generation, the four dimensions of evaluation are understandability, preference, clarity of interpretation, and relevance. Based on these definitions, participants chose which of two pairs of generated story content or guided content performed better.
Dimensions: Here provides a detailed explanation of these metrics:
  1. (1)
    Preference: Whose story/principle do you like better?
  2. (2)
    Understandability(Zhang et al., 2024): Which story would you think is in easy to understand, and using correct grammar and vocabulary.
  3. (3)
    Articulately explain: Who explains the key values more clearly and concretely for children?
  4. (4)
    Attraction: Whose story sounds more creative, contains more twists and turns and more vivid details?
  5. (5)
    Story Relevancy: Which explanations are more in line with the story’s core educational goals?
Figure 6. Story content evaluation results
Result: As the figure6 shows, GPT-4 is very close to human-generated content in terms of intelligibility. The content generated by GPT-4 is superior to that of humans in terms of logical rationality, attractiveness, and preference. ChatGPT-4’s attractiveness and explanatory clarity after linking stories to new plot creation may be due to the fact that the task takes advantage of the large language model’s ability to create logical connections of language elements based on specific instructions, and that the logical links made by the large language model are more syntactically tight. Human-generated story content is close to ChatGPT4 in terms of comprehensibility and relevance. Subsequent discussions with the evaluators indicated that the lack of readability of AI-generated stories mainly stems from the fact that the text generated by GPT-4, while consistent in language structure, sometimes lacks internal logical consistency.This observation is consistent with existing research that highlights inherent limitations(Shi et al., 2023; Lovato and Piper, 2019).
4.3.3. Evaluation of Questions generation
In the problem generation evaluation module, the plot problem and the reason problem have three similar evaluation dimensions. For plot problems, the three dimensions are: understandability, story relevance, and inspiration. The question of reason has the same first two dimensions, and the third different dimension is inclusiveness. Based on these definitions, participants were asked to choose which of two pairs of plot questions or reason questions performed better.
Dimensions: Here provides a detailed explanation of these metrics.
  1. (1)
    Understandability: Questions are generated in an easy-to-understand format, using correct grammar and vocabulary.
  2. (2)
    Story Relevancy: The questions generated relate to the story plot.
  3. (3)
    Inspiration: The generated question stimulates children’s thinking and provokes them to articulate their thoughts.
  4. (4)
    Inclusiveness: The questions generated are open-ended, allowing children to answer freely.
Result: As the figure7 shows, The evaluations between the human-generated questions and GPT-4 generated questions were close, with no significant differences. This shows that the GPT-4 model’s ability to design problems in story co-creation is acceptable and close to human level. Questions generated by humans score relatively low on story relevance compared to other dimensions. In order to maintain the continuity of the story, the relevance of the problem to the story needs to be considered. This may be because for human participants with early childhood education experience, the task of designing easy questions tends to associate them with various situations in life, resulting in the question design being too divergent and not sufficiently relevant to the story. The questions generated by GPT-4 were also more enlightening and inclusive than those generated by humans, possibly because human participants sometimes created questions that were related to educational goals but unrelated to plot development. The questions generated by GPT-4 can more meaningfully collect children’s feedback on plot development and further promote the development of the story. When it comes to reasoning, human participants sometimes create questions that are related to everyday issues, but unrelated to the core content of the story itself, and can be answered with a clear ”yes” or ”no”, which also limits the inclusiveness of the question and reduces the space for the child to answer.
Figure 7. Story questions generate evaluation results
Therefore, in reference to the opinions of the evaluators, the API interface of GPT-4 is used in this study in the formal system construction, because the story text and question text generated by GPT-4 can achieve the effect of real human guides in all dimensions of quality, and has good comprehensibility.
5. Storypark Evaluation
The experimental process includes usability testing and formal experiments. Usability testing aims to validate the usability of the interface design and ensure that participants in the formal experiment can learn to use the interface. The formal experiment focuses on evaluating whether the process of co-creating stories and answering questions about the core ideas and knowledge of the story system can help children understand the key ideas and knowledge of the story. There is no overlap between the participants in usability testing and formal experiments. The participants in usability testing have a younger average age than those in formal experiments. This is because children older than 7 years old have more experience using software on mobile phones and tablets, making it necessary to test the usability of the story system for younger participants.
Therefore, during the evaluation phase, our research questions are the following three:
  • RQ1: How would Storypark affect story knowledge and point learning outcome?
  • RQ2: How would Storypark affect the learning experience?
  • RQ3: What are educators and parents perceptions towards Storypark?
5.1. Usability test
To probe the user experience and perceived usefulness of the three storytelling phases supported by Storypark, we conduct a within-subjects usability test in a children educational institution in Minhang District, Shanghai, China.To make sure that even kindergarteners can also learn to use our interface in formal experiments, we recruited four preschool children between the age of 6 and 7 years(2 females, 2 males, mean age: 6.75 (SD = 0.5)). The purpose of usability testing in this stage is not to evaluate the effectiveness of Storypark in the story creation process, but to improve the interface design of the story system through the feedback of the children. Children can provide us with valuable feedback because they learn from stories on a daily basis in training facilities.
5.1.1. Procedure
After obtaining parental consent and child assent, participants were introduced to a research assistant and began the four-part user study procedure which consisted of introduction, understanding, interaction and evaluation. The devices be used in this experiment were a iPad and a Mac. The two research assistants involved in the experiment both have experience teaching preschoolers and both have design backgrounds.
During the introductory phase, participants will be required to view the three main screens of StoryPark on a Mac computer (i.e., co-creating the plot, understanding, and migrating the creation). The research assistant first introduced the participants to the use of the buttons in the interface and informed the children about the correct order to use the dialogue buttons and the sketch buttons. The research assistant then explains the meaning of the four icons that imply the flow of the story, and which stage of story creation each corresponds to each of the four nodes.
In the comprehension phase after the introduction phase, the research assistant points to the buttons and icons in the iPad interface and asks questions to the participants. This is to see if the child can correctly say the usefulness of the two buttons in the interface, the interaction method, the order of use, and the meaning of the icon change. The research assistant assessed the participant’s understanding of the interface using the yes/no Guttman scale. When a participant’s explanation is incorrect, or the participant is unable to provide an informative answer (e.g., silence, answer I don’t know, etc.), the research assistant will again provide the correct answer about how to use the interface to help the participant move on to the next stage.
Participants are required to complete the tasks of the three stages using the provided iPad in the interactive phase. All participants in the usability testing phase will not participate in subsequent formal experiments, so the story materials used in this testing will also be used in the actual experiments. Participants were required to complete two interactive tasks concluded two storylines co-creation through voice, two storylines co-creation through canvas, two questions on the core value of the story, and one transfer creation using new story elements. The research assistant is also present, and when the child does not respond for a long time, the research assistant will provide a small amount of help promptly, but will not directly tell the child what to do next.
Participants rated the story creation system by using a 5-point sus scale (Cronbach’s Alpha was good for the total scale (=0.78)). In the scale used, all problem descriptions were changed to colloquial descriptions that were easy for children to understand. At the same time, in terms of visual presentation, the 5-point Richter scale has also been improved to be easier for children to understand. That is, 5 arranged from small to large, drawing circles from crying faces to smiling faces. The full scale can be found in Appendix B. Considering the limited literacy skills of preschoolers, the research assistant will dictate the questions one by one, show the participants a modified Richter scale on a computer screen, explain that the crying faces to the smiling faces in the small to large circles represent ”unhappy” and ”very happy”, and ask the children which circle to decide to choose for this question.
Finally, we asked the children for their opinions on StoryPark and how they felt during the use of the Storypark. They spent about 20 minutes in the study and were given children’s toys worth 20 RMB.
5.1.2. Result
Our usability testing score is 73.33 (¿65), indicating that our interface design has good system usability among children. One participant could not recall the meaning of the left-hand process hint button; therefore, in the formal experiment, we added symbols of ”light bulb” and ”question mark” to the left-hand process buttons. Additionally, we enlarged the recording and drawing buttons on the right-hand side to facilitate children’s operation.
5.2. Participants
We sought 20 child participants (11 female, 9 male (M=7.97 S=2.81) between the ages of 5 and 13 years in child training institutions and kindergartens through offline recruitment. Gender and participant imbalances we will discuss in the ”Limits” section. None of the child participants had extensive experience with AI tools. The participants were compensated with 20 yuan worth of money for the 20-minute session. In the formal experiment, we designed story materials tailored to participants of different age groups, with varying emphasis on key ideas and knowledge. Drawing from insights provided by five educators, for participants under the age of 6, the key ideas and knowledge in the stories primarily encompass specific behaviors and everyday language usage (e.g., expressing gratitude promptly after receiving help, avoiding getting wet in the rain). For participants aged 7 and above, the key ideas and knowledge include insights that require participants to contemplate and analyze the logic of character behaviors (e.g., adopting a different perspective to tackle challenges, appreciating one’s unique qualities).
5.3. Metrics
In user study, the evaluation metrics focus was on children’s learning of learning of story value and knowledge, the quality of transferring value to the new stories they created, as well as their engagement and enjoyment throughout the process. In all instances where children’s feedback was required, an experimenter orally presented the questions and anticipated verbal responses from the children. The experimenter recorded the children’s verbal feedback, and the complete experiment was recorded on video. Thus, children did not need reading or writing skills to answer questions.
5.3.1. Learning of Story Education Goal
In previous studies, few studies focused on assessing the degree to which children understood the key idea of the story. In order to assess whether participants understood the key value of the story through co-creation with the story system, and whether the co-creation with AI had any influence on the understanding of the truth, we refer to the ”Core continent Standard”(eng, 2024) and their assessment method of exercising math ability through stories. A six-question, two-stage scale was designed. The pre-test and post-test questions were the same, and both were scored on a true/false binary scale. In order to prevent children from directly repeating the answers in the pre-test in the post-test stage, the words used in the post-test questions are slightly adjusted compared with the pre-test questions. The four dimensions of the scale refer to the definition of English Language Arts in the Common Core State Standards in the United States. Children’s understanding of the story principle was assessed through knowledge, key idea, generalization and transfer.
The knowledge and core idea dimensions are scored on a binary scale: 0 for wrong and 1 for right. Here are the definitions and sample questions for these two dimensions:
  • Knowledge: Whether the child understands the specific knowledge mentioned in the story
  • Key idea: Whether the child understands the key idea expressed by the story
For the knowledge dimension, ask the children questions about the knowledge mentioned in the story. Ask the children to explain the meaning of some common sense knowledge mentioned in the storyline, such as ”What is the characteristic of the turtle’s shell”. Each story material contains 3-5 knowledge points that can be evaluated.
In the core thought dimension, children are asked to answer questions about the truth of the story. Children need to analyze why the characters in the story did what they did and how they would move the plot if they were the characters in the story. In order to distinguish between pre-test and post-test, the description of the problem was fine-tuned in the post-test (for example, the pre-test asked, ”The little turtle does not like its shell, how do you comfort it”, and the post-test changed to ”the little turtle thinks its shell is heavy, how do you comfort it”).
The generalization and migration project is open-ended, allowing children to answer freely through it. The scores match as follows: 2 means you can transfer the key idea with a new story, 1 means you can retell the original story in your own language, and 0 means you can’t complete it. Here are the definitions and sample questions for these two dimensions:
  • Generalization: Children summarize and abstract general vocabulary (such as definition of features) after understanding the truth from specific story plots.
  • Transfer dimension: Whether children can transfer the truth of the story to a new scene and explain the truth from their own perspective.
In the generalization dimension, children are asked to understand and abstract words mentioned in stories that generalize. Such as ”What is the definition of characteristics?” . For migration problems, children need to use randomly provided new story elements to tell a new story, but still convey the core truth of the story. Each story was designed with three core principles, which were randomly selected in the experiment, and the children were asked to use new elements to create a story that could express the core principles.
Based on this scoring system, we calculate the total score by adding the scores for all the items. The Cronbach’s alpha for these problems is 0.65, which means that there is a good consistency between the four dimensions, which is acceptable.
5.3.2. Engagement
We measured children’s engagement in two ways, behavioral observations and quantitative analysis. Behavioral observation was used because child subjects sometimes have difficulty expressing their thoughts as clearly as adult subjects. Therefore, behavioral experiments can also give real feedback on children’s participation in the story park. In order to understand the participation of children in the interaction with the story park, we assessed the quality of children’s answers to questions in the quantitative indicators, and the evaluation criteria were the quality of answers and the length of answers. There are three types of response quality: meta-comments that reflect the child’s thought process (e.g., ”I think so”), statements that are irrelevant or uncertain to the question (e.g., ”I don’t know” and ”I don’t want to hear it”), and substantive responses that include meaningful details that move the story forward.
In behavioral observation, the experimenter will directly observe the following four behaviors of the participants:
  • Whether the expressions and movements are happy during the entire process of participating in the experiment;
  • Whether the tone of voice is positive when interacting with the Storypark;
  • Whether the experiment requires repeated use of certain functions of the Storypark;
  • After the experiment, are you willing to discuss the story park with the experimenter?
5.3.3. Enjoyment
To measure children’s enjoyment in interacting with artificial intelligence or humans, we adapted a questionnaire consisting of 4 items based on the prior work of Waytz(Waytz et al., 2014). These questions are as follows:
  • Are you happy when you tell a story with Storypark?
  • Do you feel comfortable telling a story with Storypark?
  • Do you want a reading companion like Storypark?
  • Do you want to tell another story with Storypark?
The experimenters verbally presented questions to the participants and guided children to indicate their level of agreement with each item using a graphical scale to express their responses. The Likert scale of five points was visualized as five circles depicting facial expressions: ”Really happy,” ”Happy,” ”Somewhat happy,” ”A little happy,” and ”Not happy at all.” Each option was associated with a circle of varying size, with circles ranging from small to large and featuring faces from crying to laughing. Researchers pointed to each circle when describing the corresponding option. The size of the circle corresponded to the degree of happiness, with the largest circle representing the greatest happiness and the smallest circle indicating the least happiness. The Cronbach’s alpha internal consistency was 0.77, which is acceptable.
5.4. Procedure and Task
After obtaining the consent of parents and children, the participants were introduced to the researchers, and the experiment was conducted in four phases: information collection, introduction, pre-test, story co-creation, narrative transfer, post-test, assessment, and interview. The devices used are ipads and Macs.
  1. (1)
    Collect information In the information-gathering phase, the children were asked some basic questions about story learning. The questions were as follows: 1. Children’s age and learning stage. 2. How often children usually listen to stories. (once a week) 3, whether the child can retell the story. 4. Does the child like to add his own ideas to the story, for example?
  2. (2)
    Introduction stage In the introduction phase, the study assistant first introduced the participants to the use of buttons in the interface, and taught the children the correct order to use the voice buttons and the canvas buttons. The research assistant then explains the meaning of the four ICONS that indicate the flow of the story, and which stage of story creation each of the four nodes corresponds to. After the research assistant asked simple questions to ensure that all children present understood the correct use of the interface, the pre-test session was conducted.
  3. (3)
    Pre-test The pre-test stage aims to test children’s understanding of the story knowledge and story truth before they interact with Storypark. Do the test by answering questions. Take the turtle story as an example. Table 5-5 lists the questions.
  4. (4)
    Story co-creation In the story creation phase, the researchers chose a story for the children according to their age group. The children need to complete four rounds of dialogue and finally create a story in cooperation with Storypark.
  5. (5)
    Narrative transfer In the narrative transfer stage, tortoise, firefly, cat, fish and other elements are given, and the task of children is to use the elements in the picture to tell a story and express the core principles of the story learned in the ”story co-creation” stage. For example, in the story co-creation stage, children create based on the story material ”the story of the turtle”, while in the narrative transfer stage, the tasks can be randomly divided into the following three: tell a story with the elements in the picture to express ”cherish your own characteristics”; Using the elements in the picture, tell a story to express ”don’t discriminate against others”; Using the elements in the picture, tell a story about ”finding the good in your friends”.
  6. (6)
    Post-test The purpose of the post-test stage is to test whether there is any change in the understanding and mastery of the story knowledge and the story truth after the children Storypark completes the co-creation of the story. Take the turtle story as an example. Table 1 lists the questions.
Table 1. List of pre and post test questions
Evaluation dimension Problem list(Correctly marked as 1 point, incorrectly marked as 0 points)
Knowledge 1. Is the turtle’s shell light or heavy?
2. From the following words, please choose the characteristics of the turtle shell: light/soft/heavy.
Key idea 3. Why did the little turtle dislike his shell at first and then become happy again?
4. If the baby turtle is sad because its shell is heavy, how do you comfort it?
Generalization 5. What does it mean to cherish?
6. What is the definition of characteristics?
Evaluation dimension Problem list(This dimension is scored 0 points in the pre-test, and 0, 1, and 2 points in the post-test respectively)
Transfer 7. Use the elements in the picture (turtle, firefly, cat, fish, bird’s nest) to tell a story and express ”cherish your own characteristics”.
8. Using the elements in the picture (turtle, firefly, cat, fish, bird’s nest), tell a story about ”finding the good in your friends”.
6. Result
6.1. RQ1: Impact on learning outcome
In order to calculate the score of this scale, we recruited 3 college students with children’s education experience to evaluate the performance of children in the experiment according to the definition of the above indicators. The data in this experiment did not meet the normal distribution, so the comparison of pre and post test data was conducted using paired sample Wilcoxon signed rank test.
Figure 8. Pre-Post Test Among four Dimensions
As can be seen from the boxplot8, the median of all indicators has increased, and the pre-test and post-test scores of core ideas and generalization have increased. The score of the four indicators of learning effect is calculated in the total score, and the median of the total score is improved, which proves that Storypark is helpful to story learning.
Table 2. Wilcoxon signed rank test for paired samples of different dimensions
Pre-test and post-test Pair Comparison P value Cohen’s d value
Question1 of Knowledge 0.025** 0.796
Question2 of Knowledge 0.157 0.459
Total score of Knowledge 0.180 0.392
Question3 of Key idea 0.001*** 1.577
Question4 of Key idea 0.000*** 1.671
Total score of Key idea 0.000*** 2.048
Question5 of Generalization 0.001*** 1.524
Question6 of Generalization 0.003*** 1.247
Total score of Generalization 0.000*** 1.949
Transfer 0.001*** 1.798
As shown in the table2, among the four dimensions, ”key idea”, ”generalization” and ”transfer” all have significant differences. The ”knowledge” dimension of problem 1 in the pre-test and post-test p values. 0.05, there was a significant difference, but question 2 of knowledge dimension and the total score of knowledge dimension were P-values in the pre-test and post-test. 0.05, no significant difference. Because the story elements prepared this time are common elements of animals, food and plants in life, and all participants have background knowledge of these elements in life, Storypark story system has little impact on this dimension.
”Core Ideas” all questions and total P-values are ¡0.05, which means that there is a significant difference between all of them, which proves that the learning process of co-creating stories with Storypark can significantly help children understand the core truth conveyed in the stories.
”Generalization” all questions and total scores show significant differences in the comparison before and after the test, which proves that Storypark can not only promote children’s understanding of the core truth conveyed in the story, but also help children abstract and summarize the truth based on specific plots in the story into the definition of some general words (such as: From the turtle shell to the generalization of the story to explain the definition of treasure, characteristics, with their own language to explain ”everyone should cherish their own characteristics of the truth”.
In the ”Transfer” dimension, 8 subjects were able to successfully complete the task of ”creating a new story that still expresses the same core idea as the original story” immediately after listening to the story. Although the closeness of the connection between the story and the core idea created by them was different, it meant that after co-creation with the AI story system, There are many children who can quickly transfer the theory to their own construction of a new scenario, and make reasonable use of their other background knowledge. Nine subjects were able to immediately retell the original story to convey the newly learned truth, while only three 5- and 6-year-old subjects were unable to construct their own story.
The ”Knowledge” dimension had no significant difference between the pre-test and the post-test, because the story elements prepared this time were common elements of animals, food and plants, and all participants already had background knowledge of these elements in their lives, so the AI story system had little impact on this dimension.
6.2. RQ2: Impact on learning experience
6.2.1. Engagement
Behavioral observation: It has been observed from children’s behavior that children generally show a positive attitude toward co-creation with story parks(Figure9)
Figure 9. Results of Children’s Engagement in the System
Among the 20 participants, 16 exhibited happy expressions, actions, and positive tones. However, 4 participants displayed confusion and perplexity, which seemed to correlate with the frequency of storytelling and story creation at home. These 4 participants reported listening to stories at home less than three times a week. Following the experiment, 8 participants actively engaged in discussions with the experimenters regarding Storypark’s functionalities. Some child participants even volunteered suggestions for Storypark, which will be addressed in RQ3. The specific experimental data are shown in the figure9. During the whole process of participating in the experiment, whether the expression and movement are happy or not (16:4); Whether participants interact with the story park in a positive tone (16:4); Whether participants were asked to use certain features of the story park repeatedly (12:8); At the end of the experiment, the participants were asked whether they would like to continue discussing the story park with the experimenter (8:12).
Additionally, 12 participants expressed a desire to use Storypark repeatedly. The functionalities most frequently requested for repeated use, ranked in descending order of frequency, were the drawing feature during plot creation (10 times), regenerating expanded text during plot creation (8 times), and drawing feature in the transfer section (5 times). There was some overlap in the number of participants making requests, with some children suggesting the repeated use of more than one feature. The “drawing-completion feature was the most frequently requested for repeated use, likely because the act of drawing with a stylus is more concrete and greatly enhances children’s sense of participation and creativity, with lower difficulty compared to the final ”Transfer creation” feature.
Quantitative analysis: Among the quantitative indicators, all subjects were able to give answers. Among all the recorded answers, 30.6% were meta-comments reflecting the thinking process, and 29.22% contained substantive statements with explanatory significance. The remaining answers that were either irrelevant to the question, or uncertain, or refused to complete the interactive task accounted for 9.8%, and the short answers of less than five words accounted for 39.2%, the highest proportion, possibly because many children answered the plot questions directly with a single word. Therefore, in terms of the combination of quantitative indicators and behavioral observation, most of the subjects actively participated in the process of co-creating stories and interacting with Storypark.
In this user experiment, the age span and study period of the subjects are large. The school section includes middle class, large class, lower grade and upper grade of primary school. Such a range of subjects can explore the adaptation degree and effect of children of different ages to the AI-assisted story co-creation system. In the experiment, it is observed that different ages have different adaptability to the story system and different willingness to use it. Based on the participation index, combined with the learning effect index and qualitative interview, it is found that the children aged 6-8 years old, in the upper class and the lower grade of primary school have the strongest adaptability to the story system. Younger children can only give short answers and rely more on the story system to help them develop, while children aged 6-8 can give rich answers and perform well in the ”migration” task, and have a strong willingness to continue working with the story system to complete the story. This means that the difficulty of the current design of the story system may be more suitable for children with certain language ability and knowledge level, and they can obtain a sense of accomplishment in the process of co-creating stories.
6.2.2. Enjoyment
Figure 10. Results of Children’s Enjoyment in the System
The figure10 shows the children’s responses to four questions (Q1 to Q4) about the pleasantness of the system. Each bar represents the percentage of responses ranging from ”strongly disagree” to ”strongly agree.” For Q1, a majority (84%) strongly agreed. Feedback from Q2 was more balanced, with 42% agreeing. Q3 shows that a majority of 21% neither agree nor disagree (gray) and 42% agree. 57% of Q4 respondents agreed, with the rest of the responses spread across other categories. Red, orange, gray, light blue, and dark blue represent different levels of disagreement and agreement, respectively.
It can be seen that children have an overall positive perception of learning story content through collaborative storytelling. There is a general high level of engagement and satisfaction: children’s willingness and satisfaction in using the AI story assistant were assessed using a four-dimensional scale ranging from 1 to 5. Among the 20 participants, 16 gave a rating of 5 for all four dimensions on the scale. Children expressed high satisfaction with the combination of voice and visual modalities. Two participants, aged 5-6, accurately articulated the value of the AI story assistant without guidance from experimenters: ”It can save my parents’ energy because it is a machine and can tell me new stories unlimited times” and ”My dad is very busy with work, he only tells me once, but AI can tell me many times”.
6.3. RQ3: Educator perceptions towards Storypark
We recruited 9 adult users, including the parents of 3 children who participated in the experiment, 3 institutional teachers who participated in the field user experiment, and 3 preschool teachers interviewed through online conference. They all tried out my AI story prototype. We conducted semi-structured interviews with parents, educators, and children to observe their attitudes towards the Storypark.
6.3.1. Storypark is an effective educational tool
Participants agreed that participating in story creation was conducive to children’s understanding of the core ideas of the story. One parent of the subjects mentioned that ”children guide the development of the story at the same time, it is a process of constantly understanding the story background and thinking about the direction of the story plot”, and he believed that this deep participation can help children understand the story more deeply. Another parent mentioned.
  • ”I think it’s good to do it in the same way that when we play RPGS (Role-Playing games), we have some plot options. When the user chooses this plot, the story goes in this direction. It makes the player feel very involved and immersed.”
One parent of the subject expressed understanding and recognition of the design of the two problem points,
  • ”When listening to the story background, he may still listen to the recording as before.” When he encounters the first problem, he will think back to what happened before. I think from his first question to his second question, he will listen or watch more carefully than he did before. Because he knew there would be a second question, and he wanted to shape the story. For example, he continues the story entirely on his own imagination, or he wants to control the development of the story according to his own preferences, so he will listen more carefully.”
6.3.2. Participating in a story brings ownership
Children become more self-aware and care more about stories that they can participate in deeply, thinking that this is ”my story.” This process allows the child to understand the truth more deeply. Children like repeated feedback, changing, rather than one-way output, the AI assistant will each time fine-tune its own extension to the child’s current feedback. The addition of visual creation can effectively attract children. At present, most picture books and story recordings are one-way or have different focuses, and few combine the two modes of speech and vision into a co-creation story system. In the current way parents tell stories and listen to recorded stories, children first hear the full plot of a story. AI Story assistants help children create storylines that add their own ideas from the start. This process of participating in the creation can help children focus their attention, have the cognition of ”this is my story”, and then have a deeper impression on the knowledge points in the story.
A parent of a 9-year-old child expressed recognition of the assessment significance of this link,
  • ”I think this link is quite meaningful. Children can organize stories in their own language, exercise their expression skills, use their imagination, and test Her ability to weave stories around core ideas is very important.”
6.3.3. Repeatedly explore the myriad possibilities in a story
Parents and education experts believe that Storypark has this significant advantage: artificial intelligence can unlimitedly embed children’s imagination into story development, and accommodate and expand children’s imagination. One parent mentioned that ”when I was a child, I probably read a lot of picture books, but the interaction was just talking and asking one or two questions.” Another parent said that he believed that his own thinking would limit his children’s imagination. In this regard, Big Language Models will perform better than humans:
  • ”Later, when she grew up, my daughter would read stories by herself and rarely read stories with us, because I felt that the child would think that we might hinder some of her understanding or some of her ideas. Because now When she grows up, she has her own thoughts. Just like when she encounters a fork in the story, she can develop from direction A or direction B. Her ideas may conflict with ours, so AI. You can accompany her to explore on her own, and the AI can completely extend the story according to her ideas.”
Children create directly without knowing the complete plot of the story, which helps them concentrate on thinking about the development of the plot, and is conducive to cultivating children’s divergent thinking and imagination. The advantage of the story system compared to human guides is that infinitely embedding children’s imagination into stories and continuously supplementing subsequent plots requires a high level of background knowledge on the part of the educator. In the classroom, teachers cannot help every child in the class generate their own stories, so they can only have a fixed story template. AI is a good creative assistant.
Both preschool teachers mentioned that adults’ thinking has accepted the teachings of social conventions and has been solidified to a certain extent, making it easy to tell stories to children with strong guidance. Because parents are so good at summarizing the truth in advance, sometimes their interpretations will constrain their children’s thinking. Some parents believe that there is only one answer to story questions and storyline interpretation, which will also affect the development of children’s divergent thinking and imagination. For parents, the process of telling the same story over and over again and replacing story elements can be tedious. Storypark can replace the process of exhausting story possibilities for children. A preschool teacher mentioned:
  • ”When the mother tells a story, she may only guide one route, and the children will only know this one route after listening to the story. AI can help children think about the story from many different angles.”
7. Discussion and future work
7.1. Experiments limitations
The participants in our study participated in only one session of AI collaborative storytelling. Such short-term visits may not fully reflect the changes that occur when children interact with the story system over a long period of time. Over time, children will become more and more accustomed to co-creating with the story system. It may take a more adaptable and interesting storytelling experience to keep a child’s attention during storytelling. Future research should explore the changes that occur after children interact with the story system over a period of time. Perhaps by giving children more opportunities to interact with the AI story system, and observing whether they will have more interaction with the story system, as well as changing some of their learning habits and thinking habits.
In addition, children’s performance in achieving educational goals after interacting with the story system may be related to some of the children’s own qualities, including background knowledge already mastered, expressive skills, introversion or extroversion, parents’ occupation, and story listening habits at home. In this experiment, for example, two six-year-old girls outperformed many older children. Specifically, they react faster in the ”migration” stage; The story is more relevant to the core values expressed in the original story; The story is created using new material rather than retelling the original story. The parents of both children are university professors, articulate and extroverted. Therefore, future story systems may be designed to provide more detailed scaffolding for introverted, slow children, especially during the final migration phase.
7.2. Challenges in the Utilization of LLMs
During the design StoryPark, we encountered three challenges while utilizing LLMs for generating questions in story continuation.
7.2.1. Impact of LLM Unpredictability Output
During the development process of interactive storytelling, we discovered that LLMs exhibit a high degree of unpredictability in their output. Slight variations in wording or syntax within the story outlines set in prompts can significantly affect the content generated by the LLMs, with each iteration of story generation producing variations. Modifications in prompt engineering cannot consistently guarantee the desired outcomes, as the interpretation of our story outlines and prompt details by LLMs remains a ”black-box” process of input and output. LLMs exhibit “dispersion” in their potential predictions for a fixed input (i.e., prompt) has also been noted by Lin et al. in their research(Lin et al., 2023). To ensure that prompt settings yield relatively stable and consistent outputs, extensive testing of story generation was conducted during development to assure the stability of the story continuations.
7.2.2. Fine-grained Control of LLMs’ Performance
LLMs demonstrate competence in understanding and executing simple directives conveyed through basic descriptive language. However, it remains a challenge when tasked with precise instructions that incorporate numerical details. Our development experience revealed that LLMs can effectively and consistently execute instructions lacking precise descriptions, such as constraints related to plot continuation or requirements tailored to specific story audience demographics. But when faced with numerically precise instructions, such as ”The story’s opening should exceed 100 words but not surpass 200 words,” LLMs frequently disregard these word count constraints, producing content shorter than 100 words. This indicates that current LLMs still have limited capabilities in understanding numerical values and reasoning, necessitating the integration of validation layers(He-Yueya et al., 2023) to ensure model outputs align with exact requirements. Feedback loops (Lee et al., 2023b) also emerges as a viable strategy.
Additionally, we encountered issues with LLMs deviating from the intended story progression, wherein the model would overlook the initial story outline and plot development constraints. To mitigate this, we employed conversational contexts to maintain a history record, and removed dialogue records from the keyword extraction phase to prevent interference with story generation, Shi also found this phenomenon(Shi et al., 2023). For longer stories, injection prompting at the end of each input has been suggested to enhance LLM performance(Liu et al., 2024).
7.2.3. Impact of LLM Hallucinations
Hallucinations in LLMs refer to instances where the generated responses possess internally flawed logic or absurd content, albeit appearing plausible(Ji et al., 2023b; Li et al., 2023). At times, LLMs may not comprehend the real-world logic of object interactions within story continuations or whether children’s feedback aligns with the question requirements, merely producing content that seems correct based on the constraints of our prompt engineering. We observed hallucinatory phenomena in our story continuations, for example, in a story about a turtle, children suggested using the turtle shell to make herbal jelly, and the LLM described the process of making herbal jelly from the turtle shell, despite the fact that turtle shells cannot be used to make herbal jelly and such a storyline cannot proceed further.
In a value-based educational system like StoryPark, such hallucinatory phenomena do not severely impact educational objectives. However, in contexts requiring scientific accuracy or logical rigor, this could lead to misunderstandings. Since the domain of mitigating hallucinations in LLMs is still in its nascent stages, future research could benefit from exploring post-processing techniques that review content generated by AI for potential hallucinatory information(Dziri et al., 2021), or employing external knowledge frameworks to guide and calibrate information generation within LLM-based conversational systems(Ren et al., 2023).
7.3. Generality for different age of Storypark
Education experts mentioned in interviews that 3-6 years old, 6-8 years old and over 8 years old are three very different stages. There are great differences in the development of narrative ability and logic ability of children in these three ages. Future research should design different story interaction modes with different difficulty levels according to children of different ages and abilities. (For example, for children with particularly strong narrative skills, co-creation questions can increase to 3-5). Improving children’s ability to understand the core values of stories through interactive stories takes time to develop habits. Future story systems should record historical data about each child’s interaction, matching varying levels of difficulty to the data the child retains in the system.
In addition, the usual learning goals are different for ages 3-6, 6-8, and 8+. Ages 3-6 usually learn specific daily habits (e.g., saying thank you regularly). Ages 6-8 begin to understand the logic and motivations of the characters’ actions in a story, and to learn general ethics and values (e.g., being helpful) from the story. Children over the age of 8 focus on learning more and more complex background knowledge from the story and analyzing more complex and multidimensional character behavior logic. The future story system should refine the generated content according to the learning content suitable for children of different ages, and generate different questions according to different learning focuses.
8. Conclusion
In this paper, we introduce Storypark, a child-story creation system that emphasizes core story values through the use of a question-scaffolding feedback framework. The storytelling system should adopt an interactive model of ”Question-Feedback-Story Generation”, presenting open-ended questions to elicit narrative contributions from children. Based on children’s contributions and the story framework, the system should expand details to support further story development. The storytelling system should augment the ability for children to create stories through voice by incorporating drawing features for visual completion, providing additional visual cues for children. The system evaluation revealed that Storypark supports creative abilities for story understanding, and is perceived as a pleasant partner accompanying the children.